Problem 2 If P = 800 N, determine the friction developed between the 60-Kg crate and the ground. The coefficient of static and dynamic friction between the crate and the ground are μs = 0.32 and μ. 0.24. respectively
Answer #1

N=P\sin \theta +mg

N=(3/5)\times 800 + 60\times 9.8 = 1068\: N

F_{s,max}=\mu_s N= 341.76

P\cos \theta = 800\times (4/5)= 640\: N

Since the horizontal force is greater than maximum static friction, the kinetic fiction comes into play.

\mathbf{F_k=\mu_k N= (0.24\times1068)\: N = 256.32\: N} (Frictional force developed)